i got 17 for my answer but it was wrong,
9/n=11/13
solving proportions

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

9/n=11/13
Multiply by 13n
117 = 11n
n = 117/11

Answer by rfer(16322)

11n=9*13
11n=117
n=117/11
n=10.6